A 30 kg block on a frictionless surface experiences two perpendicular forces in the x and y directions. If the force in the x direction is 60 N, and the force in the y direction is 150 N, find the acceleration in the x-direction, the y-direction, and the net resultant acceleration.
1
Expert's answer
2012-11-19T08:28:29-0500
acceleration in the x-directiona_x = F_x/m = (60 N)/(30 kg) = 2 m/s^2;
acceleration in the y-direction a_y = F_y/m = (150 N)/(30 kg) = 5 m/s^2;
net resultant acceleration a = sqrt((a_x)^2 + (a_y)^2) = sqrt(2^2 + 5^2) m/s^2
= sqrt(29) m/s^2 = 5.385 m/s^2.